# Contextual modulation of language comprehension in a dynamic neural model of lexical meaning

### v1

@@ 5.1 Summary & discussion
We have argued for a dynamic neural model of lexical meaning and demonstrated its behavior using the English lexical item have as a test case. The cognitive basis of the model is a continuous meaning space with two parameters: control asymmetry and connectedness. Interpretations of “possession” and “adjacency” associated with have result from the neural dynamics governing interpretation along those dimensions. The apparent discreteness of the different interpretations is epiphenomenal, resulting from the property of metastability in the neural dynamics.333The concept of metastability has previously been invoked in cognitive science most frequently to explain phenomena in visual perception, observed from both behaviorial (e.g., Chaudhuri and Glaser, 1991) and brain measures (e.g., Sterzer et al., 2009). Metastability has also been invoked to explain semantic ambiguity in language comprehension (Wildgen, 1995). Our model represents a mathematically explicit demonstration of the role of metastability in language comprehension.
@@
Simulations from the model captured known empirical effects. In particular, the interpretation evoked by English have was influenced by a preceding interpretation. An adjacency interpretation was more likely following another adjacency interpretation, and a possession interpretation was more likely following another possession interpretation. This occurred in the model because activation states persist in time, continuing to bias interpretation until they return to a resting state.444The model parameter , which controls the rate of field evolution, modulates the temporal extent of contextual bias on interpretation. Model simulations also exhibited covariation between the magnitude of contextual modulation and the strength of coupling between DNFs. Stronger coupling makes DNFs more resistant to the effects of context. In other words, there is a stronger influence of long-term knowledge, relative to immediate context, on lexical interpretation. Motivated by other observations regarding individual variation in speech behavior (described in Section 2.2), we related DNF coupling strength to the Autism-Spectrum Quotient (AQ).
Finally, model simulations generated a novel prediction: an interaction between context, acceptability, and response time. Specifically, the prediction was that in an adjacency context, more acceptable (adjacency) interpretations would be reached more quickly, while in a possession context, more acceptable interpretations would be reached more slowly. This prediction was empirically confirmed by the results of a novel experiment combining self-paced reading and acceptability judgments. Results showed an interaction between the effects of context and acceptability on response time on a trial by trial basis, such that the relationship between acceptability and response time was more negative in the adjacency context relative to the possession context. That model simulations were able to generate this empirical prediction demonstrates an important advantage of the DFT framework. In DFT, both time and cognitive features are modeled as continuous dimensions. This allows the generation of precise quantitative predictions regarding the relationship between the timecourse and outcome of cognitive processing.555In the domain of phonetics, for example, the spatial target of an articulatory movement can be related to the time it takes to generate that target (Roon and Gafos, 2016; Stern and Shaw, 2023b). As demonstrated in this study, when DNFs represent semantic dimensions, we can relate the content of a semantic interpretation to the time it takes to arrive at that interpretation. This allows us to move beyond the assumption that less acceptable sentences should always take longer to process, and thus that acceptability and response time should always be negatively correlated.
5.2 Novel predictions of our approach
@@
Our proposal that individual variation in the strength of coupling between DNFs is indexed by the AQ generates testable predictions, including in domains unrelated to linguistic meaning. For example, in an experimental task that requires learning associations between object color (one DNF) and object shape (another DNF), individuals with higher AQ scores are predicted to show greater surprisal effects (e.g., slowed response times) when encountering an object that violates the learned associations. Modeling learned dependencies between dimensions as DNF coupling may also shed light on other issues in linguistics. For example, phonological inventories (possible sounds in a language) and phonotactic constraints (possible sound sequences in a language) may be explained as the result of coupling between DNFs representing phonetic dimensions. Under this view, DNF coupling patterns are an important component of language-specific knowledge, and thus an important dimension of variation between speakers of different languages. Among speakers of the same language, processes conditioned by phonological knowledge, such as accent and perceptual illusions (Davidson and Shaw, 2012; Dupoux et al., 1999; Hallé and Best, 2007; Kabak and Idsardi, 2007), are predicted to covary with AQ scores.
5.3 Other perspectives
Before concluding, we briefly discuss the relationship of our proposal to other perspectives on lexical meaning. The proposal that lexical items map to continuous semantic spaces contrasts with models of lexical meaning based on discrete concept representations (e.g., Dell, 1986; Levelt et al., 1999; Roelofs, 1997; Stella et al., 2024). In these models, lexical meaning is formalized as a set of weighted connections between an activation node representing the lexical (e.g., CAT) and activation nodes representing discrete concept representations (e.g., ALIVE, FURRY, etc.). These models capture the intuition of the discreteness of semantic interpretations (e.g., adjacency or possession), but do not explain the evidence for a continuous substrate described in Section 1.3. In Section 2, we explained how our model fills that gap by showing how it reconciles intuitions of discreteness with evidence for continuity.
Our proposal also contrasts with distributed representations of lexical meaning, e.g., distributed parallel processing networks (e.g., Kawamoto, 1993; Rumelhart et al., 1986), or vectors describing cooccurrence patterns in written language corpora (e.g., Baayen et al., 2019; Landauer and Dumais, 1997; Mikolov et al., 2013). These models capture important facts related to the semantic similarity of lexical items, e.g., priming effects on reaction times (Kawamoto, 1993). But in contrast to our current effort, their focus is not on understanding the dimensions themselves on which lexical meanings vary. As a result, their models consist of a very large number of dimensions, each of which is not interpretable on its own. For example, in a vector-based model, the dimensionality of a word’s semantic representation is equal to the number of words in the training corpus. In a parallel distributed processing model, the dimensionality equals the number of connections between processing units. Unlike the model we proposed here, each dimension of description, on its own, is not intended to be meaningful with respect to cognition.
@@ 5.3 Other perspectives
By contrast, in our model, interpretations are modeled as peaks of activation in DNFs representing cognitively viable semantic dimensions. This model thus allows us to understand the relatedness of the various interpretations of a polysemous lexical item as proximity in the semantically parameterized space represented by DNFs. It further allows us to understand the difference between polysemy and homophony as a difference in the proximity of possible activation peaks in the space defined by the semantic DNFs to which the lexical node is coupled. Possible activation peaks are more proximal in the case of polysemy and more distal in the case of homophony. This difference arises from differences in the coupling patterns between lexical nodes and semantic DNFs. This kind of solution demonstrates an advantage of a low-dimensional “localist” framework like DFT.666This perspective also has the potential to shed light on observed differences in the timecourse of processing polysemous and homophonous words. Different interpretations of homophonous words tend to inhibit each other, leading to slowed response times, while different interpretations of polysemous words tend to prime each other, leading to faster response times (Frisson, 2015; Klepousniotou and Baum, 2007; Klepousniotou et al., 2012; MacGregor et al., 2015; Rodd et al., 2002). These facts may be explained as a result of the basic dynamics of lateral interaction within DNFs. As described in Section 2, neurons which are proximal in a DNF tend to excite each other, while those which are more distal tend to inhibit each other. These are the dynamics which allow the formation of activation peaks. Neural recordings consistent with these dynamics have been observed in rhesus monkey motor cortex (Georgopoulos et al., 1986) and cat visual cortex (Jancke et al., 1999). The difference between polyseme and homophone processing can be seen as an effect of metric feature distance on target-distractor interaction, analogous to effects observed in the domains of eye saccades (Kopecz and Schöner, 1995), manual reaching movements (Erlhagen and Schöner, 2002), and speech articulation (Tilsen, 2009).
@@ 5.4 Conclusion
Using a single theoretical framework (DFT) and mathematics of description (differential equations) allows explicit integration across the cognitive and sensory-motor domains. Previous DFT modeling work has linked neural representations of conceptual structure (Jackendoff, 2002) with visual perception (Grieben and Schöner, 2022) and visual search behavior (Sabinasz et al., 2023). Explicit coupling between abstract cognitive processes and sensation/movement has been termed “grounding” of cognition (e.g., Sabinasz et al., 2023; Sabinasz and Schöner, 2023). As described in Section 1.4, existing DFT models of speech and language have focused on the sensory-motor domain, with DNFs representing phonetic dimensions of articulatory movement and auditory perception. In this paper we have extended DFT to linguistic meaning through the lens of lexical polysemy, a cognitive domain that appears relatively distinct from sensation and behavior. Our use of the same theoretical framework and mathematics as these previous models paves the way for explicitly coupling the cognitive and sensory-motor aspects of language. This would represent a significant step towards a complete neurocognitive model of language from meaning to form: a grounded model of linguistic cognition.
@@ 6 Data availability
Data (experimental and simulated) and scripts for analysis and simulation are available on OSF at https://osf.io/3bue9/

### v2

@@ 4.1 Summary & discussion
We have argued for a dynamic neural model of lexical meaning and demonstrated its behavior using the English lexical item have as a test case. The cognitive basis of the neural model is a continuous meaning space with two parameters: control asymmetry and connectedness. Interpretations of “possession” and “adjacency” associated with have result from the neural dynamics governing interpretation along those dimensions. The apparent discreteness of the different readings results from the structure imposed by node-field coupling and field-field coupling, along with the property of nonlinearity in the neural dynamics.
@@
Simulations from the model captured known empirical effects. In particular, the specific reading evoked by English have was influenced by preceding context. While adjacency readings were overall more likely than possession readings (because the two nouns in the target sentence were unrelated and both inanimate), the likelihood of a possession reading was increased following a possession context, analogous to syntactic priming effects in comprehension (e.g., Arai et al., 2007; Ledoux et al., 2007; Tooley and Traxler, 2010). Moreover, adjacency readings were generally reached more quickly than possession readings, as reflected by the main effect of context on response time. We also related the simulation results to acceptability judgments through a generalizable measure (Eq. 7), deriving the empirical effect of context on acceptability (Zhang et al., 2018). The effects of the context sentence on target sentence interpretation occurred in the model because activation states persist in time, continuing to bias interpretation until they return to a resting state.111111The model parameter , which controls the rate of field evolution, modulates the temporal extent of contextual bias on interpretation.
@@
Model simulations also exhibited covariation between the magnitude of contextual modulation and the strength of coupling between DNFs. Stronger coupling makes DNFs more resistant to the effects of context. In other words, there is a stronger influence of long-term knowledge, relative to immediate context, on lexical interpretation. Motivated by previously reported observations regarding individual variation in speech behavior (described in Section 2.2), we related DNF coupling strength to the Autism-Spectrum Quotient (AQ). Our proposal that individual variation in the strength of coupling between DNFs is indexed by the AQ generates testable predictions, including in domains unrelated to linguistic meaning. For example, in an experimental task that requires learning associations between object color (one DNF) and object shape (another DNF), individuals with higher AQ scores are predicted to show greater surprisal effects (e.g., slowed response times) when encountering an object that violates the learned associations. Modeling learned dependencies between dimensions as DNF coupling may also shed light on other issues in linguistics. For example, phonological inventories (possible sounds in a language) and phonotactic constraints (possible sound sequences in a language) may be explained as the result of coupling between DNFs representing phonetic dimensions. Under this view, DNF coupling patterns are an important component of language-specific knowledge, and thus an important dimension of variation between speakers of different languages. Among speakers of the same language, processes conditioned by phonological knowledge, such as accent and perceptual illusions (Davidson and Shaw, 2012; Dupoux et al., 1999; Hallé and Best, 2007; Kabak and Idsardi, 2007), are predicted to covary with AQ scores.
Finally, model simulations predicted an interaction between context and acceptability in predicting response time (see Figure 14). This prediction has not previously been reported or empirically tested (although other models may generate the same prediction under certain assumptions; see Section 4.2.2). Our model generated this prediction in the following way. The adjacency context sentence reinforced the overall bias for an adjacency interpretation of the target sentence, leading to a negative correlation between acceptability and response time. The possession context sentence introduced a bias towards a possession interpretation of the target sentence, competing with the overall bias towards an adjacency interpretation. This led to a positive correlation between acceptability and response time in the possession context, but of a weaker magnitude than the negative correlation observed in the adjacency context. This prediction was partially confirmed by the results of a behavioral experiment combining self-paced reading and acceptability judgments. Specifically, results showed an interaction between context and acceptability in predicting reading time on a trial by trial basis, such that the relationship between acceptability and reading time was more negative in the adjacency context relative to the possession context (Figure 18). Also as predicted, there was a significant negative main effect of acceptability on reading time, and a negative correlation between reading time and acceptability in the adjacency context.
However, the predicted positive correlation between reading time and acceptability in the possession context was not observed. We attribute this discrepancy to a bias introduced by the experimental task. While our simulated acceptability measure is derived only from the interpretation of the target sentence (which is still, of course, influenced by the context sentence), participants were instructed to rate the acceptability of both sentences together. The perceived (in)congruity of the context and target sentences may have increased the magnitude of the negative correlation between acceptability and reading time across both contexts, eliminating the predicted (but already weak) positive correlation in the possession context, since this potential factor was not incorporated in our model simulations. This said, the key finding, predicted by the model, remains that context modulates the relationship between acceptability and reading time. We leave for future work the more detailed neural modeling of acceptability judgments (see Section 2.3) that this finding motivates.
4.2 Broader theoretical context
4.2.1 Dynamical systems models of sentence processing
Our model is consistent with a tradition of applying dynamical systems theory to human language, including the DFT work cited in Section 1.5, but also work that predates DFT or does not rely on the specific neural mechanisms of DFT. Poston (1987) and Wildgen (1995), for example, proposed that lexical semantic ambiguities are the result of multistable energy landscapes in the dynamics of cognitive states defined by continuous semantic dimensions. While remaining agnostic about the neural mechanisms that give rise to the multistabilities, they rely on the same concepts that we rely on here: continuous semantic dimensions, apparent categoricity through nonlinear dynamics, and influence of context via persistence through time of cognitive states. Our model can be seen as a more detailed extension of this work, incorporating the specific neural mechanisms from DFT and thus deriving a range of specific quantitative predictions.
Related work has applied dynamical systems ideas to language processing through the lens of connectionist or parallel distributed processing models (e.g., Kawamoto, 1993; Tabor et al., 1997; Tabor and Tanenhaus, 1999). The input units of these models are discrete localist representations, in contrast to the continuous representations in our model. In Kawamoto (1993), input units represent letters, phonemes, parts of speech, and dummy semantic variables which capture semantic similarities between lexical items. In Tabor et al. (1997), input units represent words. However, after training, network parameters form a high-dimensional continuous similarity space, between lexical items in Kawamoto (1993), and between grammatical classes in Tabor et al. (1997).
These models are similar to ours in a variety of ways: sentence processing corresponds to a trajectory through a continuous state space; linguistic interpretations correspond to attractors in the state space; and response time corresponds to the temporal duration from onset of stimulus to settling at an attractor. However, there are also crucial differences. First, the state spaces in Kawamoto (1993) and Tabor et al. (1997) are only continuous in activation (and related variables like connection weights and biases). The units over which these continuous variables are defined are themselves discrete, as described above. In our model, both activation and the units of representation over which activation is defined are continuous. Relatedly, the dimensions of the state space in our model are interpretable, and much fewer than those of Kawamoto (1993) and Tabor et al. (1997).
Moreover, the processing dynamics of our model are independently motivated. This requires less stipulation in order to link model behavior to human behavioral data. By contrast, Tabor et al. (1997), for example, specifies a mechanism of gravitational attraction to locations in state space in order to derive response time predictions. This mechanism is separate from the processing network itself, and is not related to any specific neural mechanism or principle. In our model, like that of Kawamoto (1993), attractors in state space arise from the dynamics of the processing model itself, and so no separate mechanism has to be stipulated in order to link model behavior to human response times. Our model represents a further improvement over Kawamoto (1993) in that connections between units are constrained by neural principles (in particular, local excitation and distal inhibition), rather than being free to vary during training.
An advantage of connectionist models is that the structure of the state space (representing lexical items in Kawamoto (1993) and syntactic classes in Tabor et al. (1997)) emerges through learning. We do not explicitly model the emergence of the state space structure; rather, we stipulate the structure (node-field coupling and field-field coupling) based on prior empirical evidence. This does not represent a commitment on our part that this structure does not ultimately emerge through learning; differences between languages necessitate the conclusion that some coupling relationships are formed during learning. However, it is possible that other coupling relations are innate, as suggested by the cross-linguistic similarities described in Section 1.2. By stipulating the parameters of our model, we sidestep the role of learning in this paper, focusing instead on processing in the adult system. This said, the tools exist in the DFT framework for modeling parameter learning over developmental timescales (e.g., Bhat et al., 2022). Applying these tools to the learning of lexical polysemy would be a fruitful area for future work.
4.2.2 Bayesian models of sentence processing
Models of sentence comprehension based on Bayesian inference, like surprisal theory (Levy, 2008), utilize probability distributions over possible representations; the distributions update in response to linguistic input. By stipulating that the time to comprehend a word is proportional to the Kullback-Leibler divergence between the probability distributions before and after encountering the word, surprisal theory is able to predict a variety of subtle empirical patterns in reading times. In Levy (2008), probability distributions are defined over syntactic structures, which are inherently discrete. Yet, it should be straightforward to extend this framework to lexical semantics: probability distributions would be defined over a continuous space of possible lexical semantic interpretations, rather than over a set of discrete syntactic structures. The probability distributions in Levy (2008) can be seen as approximations of the neural activation distributions in our model, since more active representations are more likely to be selected. In fact, Levy (2008) speculates that the probability distributions in his model likely arise from neural activation distributions (p. 1135, footnote 8).
Both neural activation distributions and probability distributions are dynamic, since they change in response to linguistic input. A crucial difference, though, is that the neural activation distributions in our model, and not the probability distributions in Levy (2008), have internal dynamics. They change not only in response to linguistic input, but also in response to their own internal state, via within-field and cross-field interaction. This makes our model more complex than that of Levy (2008); it is worth spelling out what justifies this additional complexity.
The first advantage of internal neural dynamics over internally-static probability distributions is that the former offers a built-in mechanism for interpretation selection: activation peak stabilization. No such mechanism exists in Levy (2008); it would have to be stipulated. A reasonable stipulation would be that interpretation selection is a sampling of the probability distribution at some timestep of processing. This stipulation correctly predicts the observed effect of context on acceptability: the selected interpretation of the target sentence would be probabilistically biased by the context sentence. Since response times are related to change in probability distribution, a further stipulation is required to predict the interaction between context, acceptability, and response time: that the selection process brings the probability of the selected interpretation to 1, and that of all other interpretations to 0. Under this set of stipulations, the Bayesian inference model could be made to generate the same predictions as the DFT model. The internal neural dynamics of DFT, while apparently increasing model complexity relative to a Bayesian model, obviate the need for such stipulations.
@@ 4.2.2 Bayesian models of sentence processing
A second advantage of the internal neural dynamics of DFT is that they generate novel predictions, e.g., by relating the interaction of representations to their metric distance in feature space. In general, interaction is excitatory for more similar representations, and inhibitory for more dissimilar representations. These are the dynamics which allow the stabilization of activation peaks; neural recordings consistent with these dynamics have been observed in cat visual cortex, for example (Jancke et al., 1999). With respect to language, it has been observed that different interpretations of homophonous words—which are more distal in semantic feature space—tend to inhibit each other, leading to slowed response times, while different interpretations of polysemous words—which are more proximal in semantic feature space—tend to prime each other, leading to faster response times (Frisson, 2015; Klepousniotou and Baum, 2007; Klepousniotou et al., 2012; MacGregor et al., 2015; Rodd et al., 2002). From the current perspective, these effects can be seen as a natural consequence of the basic dynamics of lateral interaction in neural fields, analogous to effects of metric feature distance on target-distractor interaction observed in the domains of eye saccades (Kopecz and Schöner, 1995), manual reaching movements (Erlhagen and Schöner, 2002), and speech articulation (Tilsen, 2009). Careful variation of the feature distance between interpretations in priming studies could potentially be used to empirically constrain the parameters of lateral interaction which determine the radii of excitatory and inhibitory projection in neural fields governing processing on semantic dimensions.
In summary, Bayesian inference approaches like that of Levy (2008) have an advantage of simplicity. We propose that the additional complexity in a DFT-based model like ours is worthwhile since the model generates a wider variety of novel empirical predictions, and requires less stipulation to link model behavior to human behavior. More broadly, our model offers a neural process-based account of language comprehension, which is beyond the scope of a Bayesian inference model.
@@ 4.3 Conclusion
Using a single theoretical framework (DFT) and mathematics of description (differential equations) allows explicit integration across the cognitive and sensory-motor domains. Previous DFT modeling work has linked neural representations of conceptual structure (Jackendoff, 2002) with visual perception (Grieben and Schöner, 2022) and visual search behavior (Sabinasz et al., 2023). Explicit coupling between abstract cognitive processes and sensation/movement has been termed “grounding” of cognition (e.g., Sabinasz and Schöner, 2023; Sabinasz et al., 2023). As described in Section 1.5.1, existing DFT models of speech and language have focused on phonetic dimensions of articulatory movement and auditory perception, or perceptual dimensions of objects like color, size, and spatial position. In this paper we have extended DFT to linguistic semantics through the lens of lexical polysemy, a cognitive domain that appears relatively distinct from sensation and behavior. Our use of the same theoretical framework and mathematics as these previous models paves the way for explicitly coupling the cognitive and sensory-motor aspects of language. This would represent a significant step towards a complete neurocognitive model of language linking meaning and form: a grounded model of linguistic cognition.
@@ 5 Data availability
Data (experimental and simulated) and scripts for analysis and simulation are available on OSF at Stern and Piñango (2024).
@@ 6 Funding sources
Appendix A Experimental stimuli
| stimulus set | stimulus type | condition | stimulus |
| 1 | experimental | adjacency | The maple tree has a plastic box behind it and the oak tree has a skateboard that is red. |
| 1 | experimental | possession | The maple tree has a branch that is thick and the oak tree has a skateboard that is red. |
| 2 | experimental | adjacency | The streetlight has a pink scooter next to it and the traffic sign has a shrub that is short. |
| 2 | experimental | possession | The streetlight has a bulb that is bright and the traffic sign has a shrub that is short. |
| 3 | experimental | adjacency | The wooden desk has a small bottle inside of it and the metal desk has a pen that is old. |
| 3 | experimental | possession | The wooden desk has a drawer that is wide and the metal desk has a pen that is old. |
| 4 | experimental | adjacency | The brick building has a new bus behind it and the concrete building has a crate that is large. |
| 4 | experimental | possession | The brick building has a window that is broken and the concrete building has a crate that is large. |
| 5 | experimental | adjacency | The table has a large plate on top of it and the chair has a book that is brown. |
| 5 | experimental | possession | The table has a leg that is wobbly and the chair has a book that is brown. |
| 6 | experimental | adjacency | The blanket has a soft pillow underneath it and the sheet has a bottle that is plastic. |
| 6 | experimental | possession | The blanket has a pattern that is beautiful and the sheet has a bottle that is plastic. |
| 7 | experimental | adjacency | The textbook has a small mug next to it and the notebook has a stapler that is silver. |
| 7 | experimental | possession | The textbook has a page that is torn and the notebook has a stapler that is silver. |
| 8 | experimental | adjacency | The skillet has a yellow onion next to it and the pan has a glass that is tall. |
| 8 | experimental | possession | The skillet has a handle that is sturdy and the pan has a glass that is tall. |
| 9 | experimental | adjacency | The door has a green fern next to it and the mirror has a painting that is large. |
| 9 | experimental | possession | The door has a lock that is strong and the mirror has a painting that is large. |
| 10 | experimental | adjacency | The monitor has a wooden ruler behind it and the laptop has a cup that is black. |
| 10 | experimental | possession | The monitor has a screen that is spotless and the laptop has a cup that is black. |
| 11 | filler-related | adjacency | The maple tree conceals a plastic box behind it and the oak tree has a skateboard that is red. |
| 11 | filler-related | possession | The maple tree possesses a branch that is thick and the oak tree has a skateboard that is red. |
| 12 | filler-related | adjacency | The streetlight abuts a scooter that is pink and the traffic sign has a shrub that is short. |
| 12 | filler-related | possession | The streetlight possesses a bulb that is bright and the traffic sign has a shrub that is short. |
| 13 | filler-related | adjacency | The wooden desk holds a small bottle inside of it and the metal desk has a pen that is old. |
| 13 | filler-related | possession | The wooden desk possesses a drawer that is wide and the metal desk has a pen that is old. |
| 14 | filler-related | adjacency | The brick building conceals a new bus behind it and the concrete building has a crate that is large. |
| 14 | filler-related | possession | The brick building possesses a window that is broken and the concrete building has a crate that is large. |
| 15 | filler-related | adjacency | The table holds a large plate on top of it and the chair has a book that is brown. |
| 15 | filler-related | possession | The table possesses a leg that is wobbly and the chair has a book that is brown. |
| 16 | filler-related | adjacency | The blanket covers a pillow that is soft and the sheet has a bottle that is plastic. |
| 16 | filler-related | possession | The blanket possesses a pattern that is beautiful and the sheet has a bottle that is plastic. |
| 17 | filler-related | adjacency | The textbook abuts a mug that is small and the notebook has a stapler that is silver. |
| 17 | filler-related | possession | The textbook contains a page that is torn and the notebook has a stapler that is silver. |
| 18 | filler-related | adjacency | The skillet abuts an onion that is yellow and the pan has a glass that is tall. |
| 18 | filler-related | possession | The skillet possesses a handle that is sturdy and the pan has a glass that is tall. |
| 19 | filler-related | adjacency | The door abuts a fern that is green and the mirror has a painting that is large. |
| 19 | filler-related | possession | The door possesses a lock that is strong and the mirror has a painting that is large. |
| 20 | filler-related | adjacency | The monitor conceals a wooden ruler behind it and the laptop has a cup that is black. |
| 20 | filler-related | possession | The monitor possesses a screen that is spotless and the laptop has a cup that is black. |
| 21 | filler-unrelated | metonymy | A hotel guest approaches the hotel’s front desk and asks the clerk: “Has room 451 checked out yet?” |
| 21 | filler-unrelated | no metonymy | A hotel guest approaches the hotel’s front desk and asks the clerk: “Has room 451 been cleaned yet?” |
| 22 | filler-unrelated | metonymy | In a hospital, a medical assistant asks the doctor: “Bed 22’s granddaughter is here to see him. Can she go in?” |
| 22 | filler-unrelated | no metonymy | In a hospital, a medical assistant asks the doctor: “Ed Montague’s granddaughter is here to see him. Can she go in?” |
| 23 | filler-unrelated | metonymy | One waiter in a busy restaurant complains to another: “You forgot to tell the chardonnay at table 6 about the specials!” |
| 23 | filler-unrelated | no metonymy | One waiter in a busy restaurant complains to another: “You forgot to tell the customer at table 6 about the specials!” |
| 24 | filler-unrelated | metonymy | In a diner, one waitress tells another: “The ham sandwich in the corner needs another cup of coffee.” |
| 24 | filler-unrelated | no metonymy | In a diner, one waitress tells another: “The tall woman in the corner needs another cup of coffee.” |
| 25 | filler-unrelated | metonymy | At an art shop, the owner tells the frame maker: “The Monet will pick up his order on Friday.” |
| 25 | filler-unrelated | no metonymy | At an art shop, the owner tells the frame maker: “The professor will pick up his order on Friday.” |
| 26 | filler-unrelated | metonymy | The server in a sandwich shop tells the chef: “The tuna on rye wants extra mayonnaise and a side salad.” |
| 26 | filler-unrelated | no metonymy | The server in a sandwich shop tells the chef: “The tuna on rye comes with mayonnaise and a side salad.” |
| 27 | filler-unrelated | metonymy | One stylist in a hair salon says to another: “The highlights sitting in chair 3 needs a blow-dry.” |
| 27 | filler-unrelated | no metonymy | One stylist in a hair salon says to another: “The lady sitting in chair 3 needs a blow-dry.” |
| 28 | filler-unrelated | metonymy | In a crowded emergency room, a nurse tells the doctor: “Room 3’s blood pressure is very high, and he feels dizzy.” |
| 28 | filler-unrelated | no metonymy | In a crowded emergency room, a nurse tells the doctor: “Mr. Lee’s blood pressure is very high, and he feels dizzy.” |
| 29 | filler-unrelated | metonymy | In an expensive hotel, the front desk clerk tells the chef: “Room 12 ordered room service, they want the lobster.” |
| 29 | filler-unrelated | no metonymy | In an expensive hotel, the front desk clerk tells the chef: “Mr. Dell ordered room service, they want the lobster.” |
| 30 | filler-unrelated | metonymy | A waiter in a burger joint yells to the cook: “The cheeseburger at table 10 just asked for extra pickles!” |
| 30 | filler-unrelated | no metonymy | A waiter in a burger joint yells to the cook: “The cheeseburger for table 10 should come with extra pickles!” |
| 31 | filler-unrelated | metonymy | On an airplane, a flight attendant tells the pilot: “Seat 25D will not sit down, so we can’t take off yet.” |
| 31 | filler-unrelated | no metonymy | On an airplane, a flight attendant tells the pilot: “One passenger will not sit down, so we can’t take off yet.” |
| 32 | filler-unrelated | metonymy | In a seafood restaurant, a waiter tells the busboy: “The clam chowder at table 3 ordered a glass of wine.” |
| 32 | filler-unrelated | no metonymy | In a seafood restaurant, a waiter tells the busboy: “The blonde lady at table 3 ordered a glass of wine.” |
| 33 | filler-unrelated | metonymy | In a university, a professor tells her students: “Classroom 217 is still in session, so we’ll have to wait.” |
| 33 | filler-unrelated | no metonymy | In a university, a professor tells her students: “Classroom 217 is still being used, so we’ll have to wait.” |
| 34 | filler-unrelated | metonymy | In a steak house downtown, a waitress calls to the chef: “The filet mignon wants it with rice and vegetables!” |
| 34 | filler-unrelated | no metonymy | In a steak house downtown, a waitress calls to the chef: “The filet mignon comes with rice and vegetables!” |
| 35 | filler-unrelated | metonymy | One emergency room nurse says to another: “The appendicitis in room 17B says she needs another pain pill.” |
| 35 | filler-unrelated | no metonymy | One emergency room nurse says to another: “The patient in room 17B says she needs another pain pill.” |
| 36 | filler-unrelated | metonymy | A frazzled waiter forgets the orders, so when he brings the food he asks Sue: “Are you the pasta or the steak?” |
| 36 | filler-unrelated | no metonymy | A frazzled waiter forgets the orders, so when he brings the food he asks Sue: “Is yours the pasta or the steak?” |
| 37 | filler-unrelated | metonymy | Just before the concert started, the conductor tells the orchestra: “The violins have the flu and can’t play tonight.” |
| 37 | filler-unrelated | no metonymy | Just before the concert started, the conductor tells the orchestra: “The soloists have the flu and can’t play tonight.” |
| 38 | filler-unrelated | metonymy | A barista at Starbucks this morning forgets my coffee order, then he asks me: “Are you the latte or the cappuccino?” |
| 38 | filler-unrelated | no metonymy | A barista at Starbucks this morning forgets my coffee order, then he asks me: “Is yours the latte or the cappuccino?” |
| 39 | filler-unrelated | metonymy | In an expensive French restaurant, one waiter says to another: “Table 4 asked for another bottle of chardonnay.” |
| 39 | filler-unrelated | no metonymy | In an expensive French restaurant, one waiter says to another: “That couple asked for another bottle of chardonnay.” |
| 40 | filler-unrelated | metonymy | One bartender in a cocktail lounge says to another: “The mojito at table 6 has been flirting with you all night.” |
| 40 | filler-unrelated | no metonymy | One bartender in a cocktail lounge says to another: “The girl at table 6 has been flirting with you all night.” |
| 41 | practice | grammatical | It is a beautiful day so the family is relaxing outside |
| 42 | practice | grammatical | The math test is tomorrow so the student is studying. |
| 43 | practice | ungrammatical | The game is tomorrow so the player practicing is after school. |
| 44 | practice | ungrammatical | There is a new arcade so Bill play video games every day. |